24. A prismatoid has an upper base of 90%, a lower base of 2100%, an altitude of 100 cm, and a midsection of 120? What is its volume?
1
Expert's answer
2020-12-18T14:15:50-0500

The volume of prismatoid is given by this formula:



"V=\dfrac{L}{6}(A_1+4A_m+A_2)"

where

"A_1" and "A_2=" areas of parallel bases

"A_m=" area of the section midway between "A_1" and "A_2"

"L=" the altitude



"V=\dfrac{100}{6}(90+4(120)+2100)=44500(cm^3)"